OpenStudy (anonymous):

Calculus question... Am I breaking a rule by doing this? equation soon to come

OpenStudy (anonymous):

\[{dx \over dr}={dy \over dr}\]I want to cancel out the dr's to just get \[dx = dy\]

OpenStudy (anonymous):

Then I can integrate. Can I do this?

OpenStudy (accessdenied):

that looks good to me it would be equivalent to multiplying both sides by dr and integrating \( \int \dfrac{dx}{dr} dr = \int \dfrac{dy}{dr} dr \)

OpenStudy (nincompoop):

the dr is canceled automatically

OpenStudy (accessdenied):

d/dr and integral with respect to dr are just inverse operations

OpenStudy (anonymous):

Thank you